# ACTIVA: Amortized Causal Effect Estimation without Graphs via Transformer-based Variational Autoencoder

### v1

@@ 4.4 Demonstration of Capturing the Underlying Causal Model
Figure 3 provides a qualitative view of the learned models on the held-out test Beta dataset. In this example, there is no causal effect between $X$ and $Y$. The turquoise points represent samples from the observational distribution, while the blue and brown points depict samples drawn from the interventional distributions by intervening on $X$ and $Y$ (setting them to 5).
We see that conditioning alone shifts the mean of both interventional distributions, although no causal effect is present. For example, note how the distribution of $y$ (brown points) shifts when conditioned on $x$, even though, according to the ground-truth model, it should remain unchanged. In contrast, our method correctly learns the underlying causal structure, preventing these erroneous shifts. Similar results appear in additional experiments for both the Beta and Gaussian test sets, shown in Figures 4 and 5 in the Appendix.
@@ Conclusion.
These results indicate that our method successfully leverages causal information during inference. By performing a single forward pass at test time, the model estimates accurate interventional distributions that respect the true causal mechanisms governing the data-generating process, even for unseen test data.
@@ 4.5 Performance on Semi-Synthetic Data
To evaluate our method on data that more closely resembles real-world scenarios, we employ the SERGIO (Dibaeinia & Sinha, 2020) simulator for generating gene expression data. Following the setup in (Lorch et al., 2022), we create training and test sets using a fixed set of simulator parameters, as well as an out-of-distribution (OOD) evaluation dataset with a different set of parameters. In all datasets, the intervention corresponds to a gene-knockout experiment on one of the $11$ possible genes. Appendix A.2 provides the full list of hyperparameters used in this experiment.
Table 3 presents the results of our trained model and the baseline on both the test set and the OOD evaluation set, reporting the average performance metrics over the respective datasets.The results largely align with the ones found in Section 4.4, with the exception the estimated distributions now significantly differ from the ground truth. Still, the slightly higher $p$-values for ACTIVA imply that its generated samples may deviate less from the ground truth overall.
MMD WSD ERG P
test baseline .50 5.07 3.49 .0003
ACTIVA .32 6.19 2.08 .004
ood baseline .63 8.36 8.62 .001
ACTIVA .40 10.60 4.87 .004
We also observe that performance degrades on the OOD data for both ACTIVA and the baseline. However, in addition to reflecting more challenging conditions, this decrease is partly confounded by the fact that the metrics are sensitive to data scale as detailed in Appendix B. Hence, changes in scale between the test (mean value 2.16) and OOD dataset (mean value 3.71) can amplify discrepancies in the reported metrics, even without a reduction in model performance.
@@ Conclusion.
ACTIVA exhibits robust performance on unseen semi-synthetic data, confirming its applicability to more realistic domains. Furthermore, ACTIVA remains effective even under distribution shifts, highlighting its potential for handling OOD scenarios. ACTIVA’s advantage over the conditional baseline on semi-synthetic OOD data further highlights the benefits of its causal architecture in more accurately estimating interventional distributions and its potential transfer from simulated training data to real-world test data.
@@ 5 Related Work
Deep learning methods for causal inference have made progress in modeling observational, interventional, and counterfactual distributions. Some methods utilize variational autoencoders to learn latent causal representations for high-dimensional inputs, often assuming knowledge of the causal graph or learning the causal graph (Yang et al., 2021; Qi & Yu, 2023). Diffusion-based approaches train separate generative models for each variable and require graph structures upfront (Sanchez et al., 2022; Chao et al., 2023). Other methods leverage graph neural networks or normalizing flows to model causal relations but depend heavily on predefined graphs (Zečevi et al., 2021; Sánchez-Martin et al., 2022; Javaloy et al., 2023; Poinsot et al., 2024).
Joint inference methods estimate both graph structures and causal mechanisms, but typically require retraining (Lorch et al., 2021; Nishikawa-Toomey et al., 2022; Deleu et al., 2023). Other approaches focus on estimating treatment-outcome relationships without modeling the whole distribution  (Louizos et al., 2017; Vowels et al., 2021; Wu & Fukumizu, 2023), require dataset-specific retraining (Khemakhem et al., 2021; Melnychuk et al., 2023) or are limited to specific classes of causal models (Cundy et al., 2021). Some approaches focus on special cases, such as limited overlap or marginal interventional data, but cannot handle joint distributions across all variables (Vanderschueren et al., 2023; Garrido et al., 2024).
Recently,  (Mahajan et al., 2024) explored zero-shot learning for causal inference. Their approach is restricted to additive noise models and relies on knowing the underlying causal graphs during training, restricting its applicability to real-world scenarios. In contrast, our approach amortizes causal models to directly predict the joint interventional distribution for any dataset via a single forward pass. This eliminates retraining and assumptions on the causal structure, providing a scalable, flexible solution for causal inference.
@@ 6 Conclusion
In this work, we proposed ACTIVA, a novel approach for amortized causal inference that directly predicts the interventional distribution from observational data without requiring the underlying causal graph. Our transformer-based method is both permutation equivariant and permutation invariant to variable and sample ordering, respectively, making it robust and scalable across different causal models. Furthermore, the ability to perform inference in a zero-shot manner allows ACTIVA to generalize to new datasets without retraining, making it particularly promising for many real-world applications where interventional data is scarce or unavailable.
Our experimental results demonstrate that ACTIVA accurately estimates interventional distributions across synthetic and semi-synthetic datasets. The model consistently outperforms a conditional baseline, showing its ability to capture causal relationships beyond conditioning. The results confirm that ACTIVA effectively amortizes over many causal models, providing reliable inference beyond simple point estimates even for previously unseen and OOD datasets.
Regarding the limitations, the accuracy of ACTIVA is inherently tied to the representativeness of the training data. If the training distribution does not sufficiently capture the diversity of causal models encountered during inference, the model’s predictions may still be suboptimal. Additionally, while our empirical results are promising, further theoretical work is needed to better understand the generalization properties of amortized causal learning and the conditions under which the learned representations faithfully capture underlying causal mechanisms.
Nevertheless, our work highlights the significance of amortized causal inference as a scalable alternative to traditional approaches. By leveraging the expressivity of deep generative architectures, ACTIVA represents a step toward more flexible, efficient, and broadly applicable causal reasoning. Future research directions include extending the method to handle more complex scenarios and exploring connections between amortized inference and causal identifiability. We hope this work inspires further developments in rethinking causal inference in an amortized learning context with both theoretical and practical advances.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Synthetic Data (Gaussian and Beta Noise)
where $\varepsilon_{j}$ is drawn from either: A Gaussian distribution, $\mathcal{N}(0,\sigma^{2})$, or A Beta distribution, $\mathrm{Beta}(\alpha,\beta)$. For each variable $V_{i}$, we generate interventional data by applying a single-variable intervention $\mathrm{do}(V_{i}=5)$. In total, this yields one observational dataset and $d$ interventional dataset for each causal model. All synthetic data is generated with the Causal Playground library (Sauter et al., 2024a).
When the noise terms follow a Gaussian distribution, we sample $\beta_{ij}$ uniformly from $[-2,-0.5]\cup[0.5,2]$ and let $\varepsilon_{j}\sim\mathcal{N}(0,0.5^{2})$. For the Beta case, we again draw $\beta_{ij}$ in $[-2,-0.5]\cup[0.5,2]$, but the noise terms $\varepsilon_{j}$ follow $\mathrm{Beta}(\alpha,\beta)$ with $\alpha,\beta\sim\mathrm{Uniform}(0.5,2)$.
@@
  • Two Variables. We randomly generate 6000 linear models using 3 distinct causal graphs (e.g., $X\to Y$, $Y\to X$, or no edges) and sample 50 points each for the observational and interventional distributions. We split these 6000 models into 5400 for training and 600 for testing.
  • Eight Variables. We generate 30,000 linear models (with 2000 distinct graphs, each instantiated with different parameters). For each model, we sample 30 data points for the observational and for each interventional distribution. Of these 30,000 models, 27,000 are used for training and 3000 are held out for testing.
@@ A.2 Semi-Synthetic Data (SERGIO)
To evaluate on data with biological realism, we employ the SERGIO simulator (Dibaeinia & Sinha, 2020), which generates single-cell gene-expression data. Notably, we use the version of SERGIO provided in (Lorch et al., 2022), which includes functionality for performing interventions. We treat single-gene knockouts as interventions, thus applying $\mathrm{do}(V_{i}=0)$ to each gene $V_{i}$. This yields one observational dataset and 11 single-gene interventional datasets per simulated gene-regulatory network.
@@
Following the procedure in (Marbach et al., 2009; Lorch et al., 2021), we randomly sample subgraphs of known gene-regulatory networks from E. coli or S. cerevisiae, ensuring that each subgraph has 11 genes. For each subgraph, we draw model parameters (e.g., activation constants, decay rates, and noise magnitudes) from predefined ranges. We then generate observational data as well as data from each gene-knockout intervention.
@@ Appendix B Metrics
  • Energy-Based Permutation Test
Each of these distances captures a distinct notion of distributional similarity. The MMD is a kernel-based embedding measure that detects differences in distribution means in a reproducing kernel Hilbert space; the Wasserstein distance considers a geometric “transport cost” viewpoint; the energy distance focuses on pairwise distance comparisons; and the permutation test provides a statistical criterion for whether two sets of samples could plausibly come from the same underlying distribution. Taken together, these four methods give a robust picture of how well our learned distributions match the ground truth in terms of shape, location, and higher-order moments.
@@
The MMD (Gretton et al., 2012) compares the mean embeddings of two distributions, $P$ and $Q$, in a reproducing kernel Hilbert space (RKHS). Concretely, for samples $\{\mathbf{x}_{i}\}_{i=1}^{m}\sim P$ and $\{\mathbf{y}_{j}\}_{j=1}^{n}\sim Q$, MMD is computed as:
[…]
@@
whose bandwidth parameter $\sigma$ is chosen via the median heuristic (i.e., by setting the kernel scale to the median distance among all pairs of training samples). This approach is a common, adaptive way to select an appropriate kernel width without extensive hyperparameter tuning (Garreau et al., 2017).
@@
Also referred to as the Earth Mover’s Distance, the Wasserstein distance (Villani, 2009) evaluates how much “effort” is needed to transform one distribution into another. Formally, for $p\geq 1$:
[…]
@@
The energy distance (Székely & Rizzo, 2013) offers yet another perspective on distributional similarity by comparing pairwise distances:
@@ Appendix C Hyperparameters
  • Gaussian and Beta refer to the models used in the bivariate experiments described in Section 4.4.
  • Gaussian 1 and Gaussian 10 are two versions of our architecture (with single-component and 10-component Gaussian output mixtures, respectively) used in the eight-variable experiments on Gaussian data.
  • Beta 1 and Beta 10 analogously represent single-component and 10-component mixture versions for the eight-variable Beta-noise experiments (see Appendix D.1 for results).
  • SERGIO is the version of our model employed for the semi-synthetic gene-expression data of the experiment in Section 4.5).
@@
  • lr is the learning rate used for the Adam-based optimizer (Dozat, 2016).
@@
batch size heads K dropout e d c L epochs lr seed $\beta$
Gaussian 1800 6 2 0.1 16 2 10 4 20000 0.001 42 0.01
Gaussian 1 350 6 2 0.1 16 8 1 4 10000 0.001 42 0.0005
Gaussian 10 350 6 2 0.1 16 8 10 4 9650 0.001 42 0.0005
Beta 1800 6 2 0.1 16 2 10 4 20000 0.001 42 0.005
Beta 1 350 6 2 0.1 16 8 1 4 10000 0.001 42 0.0005
Beta 10 350 6 2 0.1 16 8 10 4 9650 0.001 42 0.0005
SERGIO 256 8 2 0.1 16 8 15 4 20000 0.0005 42 0.005
@@
All runs were performed on an NVIDIA A6000 GPU, except for the SERGIO model, which was trained on an NVIDIA A100. Training times vary depending on model size (d, e, c) and dataset size (batch size, epochs). Larger mixture components ($c$) and higher embedding dimensions ($e$) generally increase both training time and representational capacity. Potential NaN’s in the optimization are zeroed.
@@
Appendix D Additional Results
In this section, we provide further experimental results that complement those presented in the main paper. We first show additional quantitative comparisons for Beta-noise models with 8 variables, and then include extended visualizations of bivariate predictions for both Beta and Gaussian data.
D.1 More Results on Beta-Noise Models
Table 6 reports the inference performance of two versions of our model, Beta 1 and Beta 10, evaluated on 3000 previously unseen test causal models with 8 variables each. These models differ only in the number of Gaussian mixture components ($1$ vs. $10$) used to parameterize the decoder output. As in the main text, we compare against the conditional baseline that fits a single multivariate normal to observational data and then conditions on the intervened variables.
Table 6: Average inference performance of our trained models and the baseline on held-out test data with eight variables. Beta1 and Beta10 are the models trained on the Beta data having single component and mixture of $10$ components, respectively.
mmd wsd erg p
Baseline .94 10.90 18.52 .00
Beta 1 .59 14.85 12.44 .00
Beta 10 .59 15.41 11.53 .002
The comparison exhibits a similar pattern as the Gaussian-noise models:
  • MMD and ERG are substantially lower for both Beta 1 and Beta 10 than for the baseline, indicating that our method’s samples align more closely with the true interventional distributions in a kernel embedding and pairwise-distance sense.
  • The WSD is higher for both Beta 1 and Beta 10. We hypothesize that this is due to the model outputting broader variances in some directions, leading to higher “transport cost” under the Wasserstein metric.
  • The permutation test $p$-value remains relatively low for all methods (including ours), suggesting some residual mismatch to the true distribution.
D.2 Visualizations of Bivariate Estimates
To complement our quantitative results, Figures 4 and 5 provide qualitative illustrations of the learned interventional distributions for bivariate models with either Beta or Gaussian additive noise. Each figure consists of four rows, each corresponding to a distinct data-generating process (or parameter configuration). In each subplot:
  • The left column (“Original”) shows samples from the ground-truth observational (turquoise) and interventional distributions (blue/brown).
  • The middle column (“Baseline”) shows reconstructed samples from conditioning on the intervened variables in a single fitted MVN.
  • The right column (“Model”) shows the corresponding draws from ACTIVA’s learned interventional distributions.
In Figures 4 and 5, we observe:
  • False Shifts in the Conditional Baseline. Conditioning alone often shifts non-target variables, even when there is no causal path. This is particularly visible for certain Beta-noise cases.
  • Correct Zero-Effect Modeling. ACTIVA preserves the distribution of non-target variables unless there is a true causal relationship, avoiding unnecessary changes.
  • Increased Variance. Our model sometimes outputs broader distributions, which can manifest in higher Wasserstein distance but still yield more faithful coverage of the true data-generating process.
Overall, these additional results reinforce our main findings. ACTIVA accurately captures causal asymmetries in both Beta and Gaussian noise settings and avoids the pitfall of naively “conditioning away” correlations to produce artifacts where no causal effect exists.
(d) (a) Ground Truth
(e) (b) Conditional Baseline
(f) (c) ACTIVA
Figure 4: Samples from the observational and interventional distributions of the ground-truth model (a), the conditional baseline (b) and our model (c) for linear models with additive Beta noise. The observational distributions of the original data and the conditional baseline are turquoise. The samples from the true and estimated distributions $p(Y\mid do(X=5))$ are blue, the ones from the distribution $p(X\mid do(Y=5))$ are brown. Note that there are no samples from the observational distribution from our model, since it does not estimate observational distributions. Instead we provide the observational samples of the original data for visualization.
(a) (a) Ground Truth
(b) (b) Baseline
(c) (c) ACTIVA
Figure 5: Samples from the observational and interventional distributions of the ground-truth model (a), the conditional baseline (b) and our model (c) for linear models with additive Gaussian noise. The observational distributions of the original data and the conditional baseline are turquoise. The samples from the true and estimated distributions $p(Y\mid do(X=5))$ are blue, the ones from the distribution $p(X\mid do(Y=5))$ are brown. Note that there are no samples from the observational distribution from our model, since it does not estimate observational distributions. Instead we provide the observational samples of the original data for visualization.

### v2

@@ 6.4 Distribution Shift Estimation
Figure 2 provides a qualitative view of the learned models on the held-out test Beta dataset. In this example, there is a negative causal effect from $Y$ to $X$.
We observe that both the baseline and ACTIVA recover the mean in the causal direction (brown), although ACTIVA has higher uncertainty. More interestingly, the baseline predicts a distribution shift in the anti-causal direction (blue). In contrast, ACTIVA correctly predicts no causal effect. Furthermore, it provides a low-probability mode in the causal direction, making uncertainty about the causal model explicit.
@@ Conclusion.
These results indicate that our method successfully leverages causal information during inference. By performing a single forward pass, the model estimates accurate interventional distributions that respect the true causal mechanisms governing the data-generating process.
ACTIVA Baseline
Data MMD WSD ERG MMD WSD ERG
SERGIO 0.34 4.3 1.7 0.50 3.9 2.8
OOD 0.37 7.6 3.7 0.61 8.0 8.0
@@ 6.5 Performance on Semi-Synthetic Data
To highlight the potential application of ACTIVA to real scenarios, we employ the SERGIO (Dibaeinia and Sinha 2020) simulator for generating gene expression data. Following the setup in (Lorch et al. 2022), we create training and test sets with 8 variables, as well as an out-of-distribution (OOD) evaluation dataset with 11 variables. Appendix A.2 provides a detailed list of the used parameters.
Table 2 compares the baseline to ACTIVA trained on the SERGIO data on the test set and the OOD evaluation set, reporting the average performance metrics. The results largely align with the ones found in Section 6.4, with the exception of almost equal performance on the WSD.
We also observe that performance degrades on the OOD data for both ACTIVA and the baseline. However, in addition to reflecting more challenging conditions, this decrease is partly confounded by the fact that the metrics are sensitive to data scale as detailed in B. Hence, changes in scale between the test (mean value 2.16) and OOD dataset (mean value 3.71) can amplify discrepancies in the reported metrics, even without a reduction in model performance.
@@ Conclusion.
ACTIVA exhibits robust performance on unseen semi-synthetic data, confirming its applicability to more realistic domains. Furthermore, ACTIVA remains effective even under distribution shifts, highlighting its potential for handling OOD scenarios. It’s advantage over the conditional baseline on semi-synthetic OOD data further highlights the benefits of its causal architecture for estimating interventional distributions and its potential transfer from simulated training data to real-world test data.
@@ 7 Related Work
Deep learning approaches to causal inference have advanced modeling of observational, interventional, and counterfactual distributions, yet most demand known or learnable causal graphs and retraining for each new dataset. This holds for variational‐autoencoder–based models (Yang et al. 2021; Qi and Yu 2023), diffusion generative models (Sanchez et al. 2022; Chao et al. 2023), graph‑neural and flow‑based approaches (Zečevi et al. 2021; Sánchez-Martin, Rateike, and Valera 2022; Poinsot et al. 2024), adversarial network approaches (Rahman and Kocaoglu 2024).
Recent amortized estimators leverage prior‑fitted networks or supervised learning to predict causal effect distributions without additional training or known causal graphs (Balazadeh et al. 2025; Robertson et al. 2025; Ma et al. 2025). However, they still rely on explicit covariate specification and typically target predefined estimands. Other meta-learning frameworks are restricted to supervisedeffect estimation (Bynum et al. 2025), enable zero-shot estimation for unseen interventions  (Nilforoshan et al. 2023), model the noise as latent  (Pawlowski, Castro, and Glocker 2020; De et al. 2023) within the same causal model context, require upfront model assumptions (Löwe et al. 2022) and graphs (Mahajan et al. 2024).
Joint inference methods that jointly learn graphs and mechanisms likewise require dataset‑specific optimisation (Lorch et al. 2021; Nishikawa-Toomey et al. 2022; Deleu et al. 2023). Other approaches focus on estimating treatment-outcome relationships without modeling the whole distribution  (Louizos et al. 2017; Vowels, Cihan Camgoz, and Bowden 2021; Wu and Fukumizu 2023), require dataset-specific retraining (Khemakhem et al. 2021; Melnychuk, Frauen, and Feuerriegel 2023) or are limited to specific classes of causal models (Cundy, Grover, and Ermon 2021). Some approaches focus on special cases, such as limited overlap or marginal interventional data, but cannot handle joint distributions across all variables (Vanderschueren et al. 2023; Garrido et al. 2024).
In contrast, our method amortizes across both causal structures and mechanisms, requiring only observational data and a specified intervention to return the full joint interventional distribution in a single forward pass. It therefore removes the need for retraining, covariate decomposition, or graphical assumptions.
@@ 8 Conclusion
We introduce ACTIVA, a VAE architecture for amortized causal distribution estimation. During training, the encoder maps interventional data — conditioned on observational inputs and a query intervention — to a latent representation, and the decoder reconstructs the corresponding interventional distribution. At inference time, conditioning solely on observational data yields a prior over plausible causal models, which the decoder transforms into a mixture of model-specific interventional distributions. We provide a rigorous theoretical analysis of these latent and output spaces and empirically validate our implementation on synthetic and semi‑synthetic datasets. ACTIVA’s primary strengths include its capability for zero-shot inference using observational data alone. This work is among the first works that offer a theoretical foundation for amortized causal inference.
Future work will focus on bridging the gap between theoretical assumptions and implementation, thoroughly evaluating ACTIVA’s performance on real-world datasets to better understand the simulation-to-reality transfer, and exploring alternative deep learning architectures aligned with the developing theory of amortized causal inference. We hope this work inspires further developments in rethinking causal inference in an amortized learning context with both theoretical and practical advances.
@@ A.1 Synthetic Data (Gaussian and Beta Noise)
where $\varepsilon_{j}$ is drawn from either: A Gaussian distribution, $\mathcal{N}(0,\sigma^{2})$, or A Beta distribution, $\mathrm{Beta}(\alpha,\beta)$. For each variable $V_{i}$, we generate interventional data by applying a single-variable intervention $\mathrm{do}(V_{i}=5)$ plus some noise $\mathcal{N}(0,0.1)$ to increase numerical stability. In total, this yields one observational dataset and $d$ interventional dataset for each causal model. All synthetic data is generated with the Causal Playground library (Sauter, Acar, and Plaat 2024).
The causal graphs are generated according to the ER procedure (Erdós and Rényi 1959) where first we generate an ER graph with edge-probability $0.3$ and then randomly remove edges until the graph becomes acyclic. When the noise terms follow a Gaussian distribution, we sample $\beta_{ij}$ uniformly from $2,-0.5]\cup[0.5,2]$ and let $\varepsilon_{j}\sim\mathcal{N}(0,0.5^{2})$. For the Beta case, we again draw $\beta_{ij}$ in $2,-0.5]\cup[0.5,2]$, but the noise terms $\varepsilon_{j}$ follow $\mathrm{Beta}(\alpha,\beta)$ with $\alpha,\beta\sim\mathrm{Uniform}(0.5,2)$.
@@
  • Two Variables. We randomly generate 10000 linear models and sample 40 points each for the observational and interventional distributions. We split these 10000 models into 9000 for training and 1000 for testing.
  • Eight Variables. We generate 30000 linear models. For each model, we sample 30 data points for the observational and for each interventional distribution. Of these 30000 models, 27000 are used for training and 3000 are held out for testing.
@@ A.2 Semi-Synthetic Data (SERGIO)
To evaluate on data with biological realism, we employ the SERGIO simulator (Dibaeinia and Sinha 2020), which generates single-cell gene-expression data. Notably, we use the version of SERGIO provided in (Lorch et al. 2022), which includes functionality for performing interventions. We treat single-gene knockouts as interventions, thus applying $\mathrm{do}(V_{i}=0)$ to each gene $V_{i}$. This yields one observational dataset and 8 single-gene interventional datasets per simulated gene-regulatory network.
@@
Following the procedure in (Marbach et al. 2009; Lorch et al. 2021), we randomly sample subgraphs of known gene-regulatory networks from E. coli or S. cerevisiae, ensuring that each subgraph has 8 genes. For each subgraph, we draw model parameters (e.g., activation constants, decay rates, and noise magnitudes) from predefined ranges (Table 3. We then generate observational data as well as data from each gene-knockout intervention.
@@ Appendix B Metrics
Each of these distances captures a distinct notion of distributional similarity. The MMD is a kernel-based embedding measure that detects differences in distribution means in a reproducing kernel Hilbert space; the Wasserstein distance considers a geometric “transport cost” viewpoint; and the energy distance focuses on pairwise distance comparisons; Taken together, these three metrics give a robust picture of how well our learned distributions match the ground truth in terms of shape, location, and higher-order moments.
@@
The MMD (Gretton et al. 2012) compares the mean embeddings of two distributions, $P$ and $Q$, in a reproducing kernel Hilbert space (RKHS). Concretely, for samples $\{\mathbf{x}_{i}\}_{i=1}^{m}\sim P$ and $\{\mathbf{y}_{j}\}_{j=1}^{n}\sim Q$, MMD is computed as:
$\mathrm{MMD}^{2}(P,Q)\;=\;\mathbb{E}_{\mathbf{x},\mathbf{x}^{\prime}\sim P}[k(\mathbf{x},\mathbf{x}^{\prime})]\;+\;\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim Q}[k(\mathbf{y},\mathbf{y}^{\prime})]\;-\;2\,\mathbb{E}_{\mathbf{x}\sim P,\mathbf{y}\sim Q}[k(\mathbf{x},\mathbf{y})],$
@@
whose bandwidth parameter $\sigma$ is chosen via the median heuristic (i.e., by setting the kernel scale to the median distance among all pairs of training samples). This approach is a common, adaptive way to select an appropriate kernel width without extensive hyperparameter tuning (Garreau, Jitkrittum, and Kanagawa 2017).
@@
Also referred to as the Earth Mover’s Distance, the Wasserstein distance (Villani 2009) evaluates how much “effort” is needed to transform one distribution into another. Formally, for $p\geq 1$:
$\mathcal{W}_{p}(P,Q)\;=\;\biggl{(}\,\inf_{\gamma\,\in\,\Gamma(P,Q)}\int_{\mathcal{X}\times\mathcal{X}}d(\mathbf{x},\mathbf{y})^{p}\,d\gamma(\mathbf{x},\mathbf{y})\biggr{)}^{1/p},$
@@
The energy distance (Székely and Rizzo 2013) offers yet another perspective on distributional similarity by comparing pairwise distances:
@@ Appendix C Hyperparameters
  • Gauss 2 and Beta 2 refer to the models used in the bivariate experiments described in 6.4.
  • Gauss 8 and Beta 8 analogously represent versions for the eight-variable models.
  • SERGIO is the version of our model employed for the semi-synthetic gene-expression data of the experiment in 6.5).
@@
  • h hidden dimension of the neural networks throughout the model.
@@
  • k denotes the number of Gaussian components used in the mixture for the prior distribution.
@@
  • lr is the learning rate used for the Adam-based optimizer (Dozat 2016).
@@
  • GPU on which the model was trained.
batch size heads K dropout e h d c k L epochs lr seed $\beta$ GPU
Gauss 2 256 2 4 0.1 64 128 2 1 2 2 5000 0.001 42 0.1 A6000
Beta 2 256 2 4 0.1 64 128 2 1 2 2 5000 0.001 42 0.1 A6000
Gauss 8 256 4 4 0.1 64 512 2 1 2 4 3600 0.0005 42 0.1 A6000
Beta 8 256 4 4 0.1 64 512 2 1 2 4 3600 0.0005 42 0.1 A6000
SERGIO 256 4 4 0.1 64 256 2 1 2 4 4100 0.0005 42 0.1 A100
@@
Training times vary depending on model size (d, e, c) and dataset size (batch size, epochs). Larger mixture components ($c$) and higher embedding dimensions ($e$) generally increase both training time and representational capacity. Potential NaN’s in the optimization are zeroed. All modules and training scripts were implemented with Jax (Bradbury et al. 2018) and Flax (Heek et al. 2024). Experiment were run on the DAS6 computing cluster (Bal et al. 2016).